# Atomic forces from correlation energy functionals based on the adiabatic- connection fluctuation-dissipation theorem

### v1

@@ VI Conclusions and outlook
In this work, we have derived and implemented atomic forces within the RPA method. Forces were evaluated at full self-consistency using the OEP method, and, non-self-consistently, using density functional perturbation theory with the PBE functional. Both self-consistent and non-self-consistent forces were shown to be of high numerical quality. Similarly to EXX and hybrid functionals, a crucial extra force term remains at full self-consistency when nonlocal pseudopotentials are employed.
The RPA force implementation has been applied to molecular and solid state systems, highlighting a systematic improvement of structural and vibrational properties with respect to simpler semi-local functionals. The errors in RPA, in particular for molecules, are found mainly in bond angles and bond-stretching modes, which are not as well described as expected. On the other hand, in diamond, RPA strongly improves the zone-center optical phonon. The error, taking into account the anharmonic shift, drops from 4.3% in PBE to 0.5% in RPA. Similar improvements are found for silicon and germanium. Overall, the error on the vibrational frequencies in RPA seems smaller in solids than molecules.
Despite the current analytical force implementation being available only for RPA, the finite-difference total energy estimates made at the RPAx level show an improvement over RPA. In fact, an excellent agreement with CCSD(T) is found for all molecules. On diamond, our RPAx phonon calculation provides the best theoretical estimate obtained so far. Future works should therefore aim to implement the exact-exchange kernel contribution, to gain access to RPAx analytical forces. A non-self-consistent implementation should be sufficient, as demonstrated for RPA in this work.
Finally, the current RPA force implementation represents an important step towards the analytical calculation of second-order derivatives of the total energy with respect to atomic positions and electron-phonon couplings. Another important future direction is the evaluation of the stress tensor, which would enable full cell relaxation.
@@
In this Appendix, we present convergence studies for diamond. The $\mathbf{k}$-point convergence is presented in Fig. 8. The behavior is very similar for all methods studied, thus allowing us to extrapolate the RPA, RPAx, RPAx(1) and AC-SOSEX results with high precision. The convergence with respect to plane-wave cutoff is presented in Fig. 9. The ACFDT-based methods are seen to converge faster with respect to this parameter, as compared to PBE and hybrid functionals.
@@
In Fig. 11, we see that for the smaller values of cutoff, a discrepancy exists between the total energy finite-difference and force finite-difference estimates. This behavior disappears, however, when increasing the value of the cutoff. While RPA[PBE] is barely impacted, the PBE and PBE0 hybrid functional are. For them, a perfect agreement is only reached at very large cutoff values.

### v2

@@ VI Conclusions and outlook
In this work, we have derived and implemented atomic forces within the RPA method. Forces were evaluated at full self-consistency using the OEP method, and non-self-consistently using density functional perturbation theory with the PBE functional. Both self-consistent and non-self-consistent forces were shown to be of high numerical quality. Similarly to EXX and hybrid functionals, a crucial extra force term remains at full self-consistency when nonlocal pseudopotentials are employed.
The RPA force implementation has been applied to molecular and solid-state systems, highlighting a systematic improvement of structural and vibrational properties with respect to simpler semi-local functionals. The errors in RPA, in particular for molecules, are found mainly in bond angles and bond-stretching modes, which are not as well described as expected. On the other hand, in diamond, RPA strongly improves the zone-center optical phonon. The error, taking into account the anharmonic shift, drops from 4.3% in PBE to 0.5% in RPA. Similar improvements are found for silicon and germanium. Overall, the error on the vibrational frequencies in RPA seems smaller in solids than in molecules.
Despite the current analytical force implementation being available only for RPA, the finite-difference total energy estimates made at the RPAx level show an improvement over RPA. In fact, an excellent agreement with CCSD(T) is found for all molecules. On diamond, our RPAx phonon calculation provides the best theoretical estimate obtained so far. Future works should therefore aim to implement the exact-exchange kernel contribution to the gradient, to gain access to RPAx analytical forces. A non-self-consistent implementation should be sufficient, as demonstrated for RPA in this work.
Finally, the current RPA force implementation represents an important step towards the analytical calculation of second-order derivatives of the total energy with respect to atomic positions, and electron-phonon couplings. Another important future direction is the evaluation of the stress tensor, which would enable full cell relaxation.
@@
In this Appendix, we present convergence studies for diamond. The $\mathbf{k}$-point convergence is presented in Fig. 8. The behavior is very similar for all methods studied, thus allowing us to extrapolate the RPA, RPAx, RPAx(1), and AC-SOSEX results with high precision. The convergence with respect to the plane-wave cutoff is presented in Fig. 9. The ACFDT-based methods are seen to converge faster with respect to this parameter, as compared to PBE and hybrid functionals.
@@
In Fig. 11, we see that for the smaller values of the cutoff, a discrepancy exists between the total energy finite-difference and force finite-difference estimates. This behavior disappears, however, when increasing the value of the cutoff. While RPA[PBE] is barely impacted, the PBE and PBE0 hybrid functional are. For them, a perfect agreement is only reached at very large cutoff values.
Figure 8: Convergence of the $\Gamma$-point optical phonon frequency of diamond with respect to the size of the sampling grid. The difference in frequency evaluated at a given $\mathbf{k}$-point and $\mathbf{q}$-point grid definition is given with respect to the frequency calculated on a $\times5$$\times$5 grid for RPAx, RPAx(1), AC-SOSEX, a $\times6$$\times$6 grid for RPA, and a $\times10$$\times$10 grid for PBE, PBE0(17%), and PBE0(34%). The RPA-based methods’ curves have been shifted upwards for extrapolation. The ACFDT calculations were performed with a kinetic energy cutoff of 80 Ry, and a total of 80 eigenvalues.
Figure 9: Convergence of the $\Gamma$-point optical phonon frequency of diamond with respect to the kinetic energy cutoff. The difference in frequency evaluated for a given cutoff is given with respect to the value calculated at 200 Ry. The PBE, PBE0(17%), and PBE0(34%) calculations have been performed on a $\times8$$\times$8 $\mathbf{k}$-point grid, while the ACFDT-based methods have been performed on a $\times3$$\times$3 grid. A total number of 80 eigenvalues was used.
Figure 10: Convergence of the $\Gamma$-point optical phonon frequency of diamond with respect to the total number of eigenvalues of the response function. The difference in frequency evaluated for a given number of eigenvalues is taken with respect to the frequency calculated at 140 eigenvalues. Calculations have been performed on a $\times3$$\times$3 and a $\times4$$\times$4 $\mathbf{k}$-point grid. A plane-wave cutoff of 80 Ry was used.
Figure 11: Difference in the $\Gamma$-point optical phonon frequency estimated by finite difference of the total energy and by finite difference of forces via the Phonopy software. The calculations have all been performed considering a $\times3$$\times$3 $\mathbf{k}$-point grid, and a total number of 80 eigenvalues for RPA[PBE]. The curves have all been shifted with respect to the RPA[PBE] frequency value calculated by finite difference at 200 Ry.